# RL in Name Only? Analyzing the Structural Assumptions in RL post-training for LLMs

### v2

@@ Abstract
Reinforcement learning-based post-training of large language models (LLMs) has recently gained attention, particularly following the release of DeepSeek R1 [8], which applied GRPO for fine-tuning. Amid the growing hype around improved reasoning abilities attributed to RL post-training, we critically examine the formulation and assumptions underlying these methods. We start by highlighting the popular structural assumptions made in modeling LLM training as a Markov Decision Process (MDP), and show how they lead to a degenerate MDP that doesn’t quite need the RL/GRPO apparatus. The two critical structural assumptions include (1) making the MDP states be just a concatenation of the actions–with states becoming the context window and the actions becoming the tokens in LLMs and (2) splitting the reward of a state-action trajectory uniformly across the trajectory. Through a comprehensive analysis, we demonstrate that these simplifying assumptions make the approach effectively equivalent to an outcome-driven supervised learning. Our experiments on benchmarks including GSM8K and Countdown using Qwen-2.5 base models show that iterative supervised fine-tuning, incorporating both positive and negative samples, achieves performance comparable to GRPO-based training. We will also argue that the structural assumptions indirectly incentivize the RL to generate longer sequences of intermediate tokens–which in turn feeds into the narrative of “RL generating longer thinking traces.” While RL may well be a very useful technique for improving the reasoning abilities of LLMs, our analysis shows that the simplistic structural assumptions made in modeling the underlying MDP render the popular LLM RL frameworks and their interpretations questionable.
@@ 1 INTRODUCTION
Recent advances in Large Language Models (LLMs) have substantially improved their ability to perform complex reasoning tasks, such as mathematical problem-solving. Typically, this is achieved through post-training techniques, including Reinforcement Learning (RL) methods like Group Relative Policy Optimization (GRPO). These methods frame language modeling as a Markov Decision Process (MDP), where token generation is treated as a sequential decision-making task.
@@
In this work, we critically examine the foundations of applying RL to LLM post-training through the lens of the specific MDP formulation that has been popularly used to model LLMs as agents  [8]. Under this formulation, each state is a sequence of previously generated tokens, transitions are deterministic concatenations of tokens, rewards are assigned only at the terminal state based on external verification of correctness, and the credit assignment problem is short-circuited by effectively dividing this terminal reward equally among the tokens (actions). These structural assumptions sharply distinguish the setting from the general MDP framework typically considered in RL leading to wrong intuitions.
@@
Our contributions are twofold. Firstly, we theoretically show that under these structural assumptions, making reinforcement learning methods post-training is somewhat redundant. By decoupling and simplifying the GRPO objective for trajectories labeled as positive (correct solutions) and negative (incorrect solutions), we demonstrate that iterative supervised fine-tuning (SFT) guided by an external verifier yields comparable performance. This equivalence implies that RL might not be imparting genuinely novel reasoning capabilities to the LLM but rather leveraging latent capabilities inherent within the pretrained models.
@@
Secondly, we empirically demonstrate that the terminal reward distribution assumption, wherein the relative advantage score is uniformly allocated across all tokens in a response, directly contributes to the increasing response lengths observed throughout training. Contrary to the claims made in the DeepSeek-R1 [8] paper—which attributes increased response length at test-time to computational scaling, self-reflection, self-verification, and back-tracking—we show that the primary driver of response elongation is this uniform credit distribution. Since each token is incentivized equally without considering the "return-to-go" from any intermediate state, RL-based methods like GRPO inadvertently encourage verbose outputs.
This paper thus clarifies misconceptions regarding the role of reinforcement learning in enhancing LLM capabilities and provides a comprehensive theoretical and empirical analysis of how degenerate MDP assumptions influence model behavior during fine-tuning. Our findings suggest both considering alternate MDP formulation (eg the 2-LLM one in [15]) and revisiting traditional supervised fine-tuning methods with external verification, and highlight that simpler approaches may achieve similar, if not better, efficiency and effectiveness compared to complex RL-based post-training strategies.
@@ 2 RELATED WORK
Supervised fine-tuning (SFT) and reinforcement learning (RL) have emerged as the two main paradigms for post-training LLMs. Although early work focused heavily on SFT variants, recent efforts have shifted toward RL-inspired methods such as GRPO and its extensions.
@@
While the original SFT involved finetuning on the traces and solutions supplied externally, a line of work has explored improving language model reasoning via self-generated data, filtering data with verifier and iteratively fine-tuning on the filtered data, a process which we refer to as filtered iterative SFT in our work. Self-Taught Reasoner (STaR) [21] fine-tunes a model on its self generated data by iteratively producing rationales using greedy decoding. It retains only samples whose rationales lead to correct answers. Subsequent works have replaced greedy decoding with temperature sampling as an alternative data generation strategy. Rejection Sampling Fine-Tuning (RFT) [20] collects reasoning paths by generating multiple outputs from a supervised model and filtering them using binary feedback (accepted/rejected). The model is then fine-tuned once on the accepted samples. RAFT (Reward-ranked Fine-Tuning) [5] introduces an iterative framework consisting of three steps: (1) generate a batch of outputs from the model; (2) score the outputs using a reward function and filter high-reward samples and (3) fine-tune the model on the filtered data.
@@
Reinforced Self-Training (ReST) [7] similarly generates multiple output sequences per prompt, scores them with a reward function, and fine-tunes the model using a reward-weighted loss, rather than hard filtering. ReSTEM [14] extends this setup by fine-tuning the original base model in each iteration, instead of the latest model checkpoint. These methods collectively explore different ways to combine self-generated data, filtering, and reward signals for language model fine-tuning.
@@
The recent and growing work has focused on variations to GRPO Framework. GRPO [12] introduces a group-relative advantages, avoiding state-based value estimation by directly assigning advantages based on relative scores within a group. Several follow-up works build on or modify GRPO in different ways. Dr-GRPO [9] removes standard deviation normalization and sample-length scaling from the GRPO loss, and points out that the objective may introduce a bias toward longer responses. GRPO-Lead [22] extends GRPO by adding a length-dependent reward to encourage concise outputs, an explicit penalty to reduce low-precision responses, and an advantage reweighting strategy to handle limited rollout data.
@@
In parallel, DAPO [18] introduces a different set of techniques for long chain-of-thought settings. It uses token-level policy gradient loss and includes four key components: Clip-Higher to maintain policy diversity, Dynamic Sampling to improve training stability, Overlong Reward Shaping to reduce noise from long outputs, and the removal of flat-reward samples from training. VAPO [19] builds on DAPO by introducing value model pretraining, a decoupled and length-adaptive GAE, and group sampling to better handle sparse rewards and variable-length responses.
@@
In this work, We argue that GRPO and its variants, despite their RL framing, rely on structural simplifications such as degenerate MDPs and uniform token-level reward assignment, which reduce the actual learning process to iterative filtered supervised fine-tuning. A Very recent work, RAFT++ [16] also extended RAFT by incorporating importance sampling and clipping and shows comparable performance as GRPO.
Further, there have been claims that the longer responses reflect greater reasoning, with increased length often linked to higher accuracy and emergent capabilities. In response, recent works aim to reduce output length while maintaining accuracy, particularly with sample level loss in DAPO[18] and Dr. GRPO [9], length-dependent penalties in [19, 1, 2]. However, we show this is more likely a side effect of the uniform advantage distribution followed with length scaling rather than true emergent reasoning. Further, [6] shows theoretical analysis that generation of longer responses stems from RL-based optimization during training.
@@ 3.1 Standard MDP Formulation
3.2 LLM-MDP
Let’s now consider the popular MDP model used for LLMs introduced in DeepSeek-R1 paper [8]. In this scenario, the LLM acts as an agent interacting with an environment whose state space is the set of all possible token sequences up to a maximum length $H$, defined as $S=\bigcup_{k=0}^{H}\mathcal{V}^{k}.$ with $\mathcal{V}$ representing a finite vocabulary of tokens. The action space $A$ is precisely this vocabulary $\mathcal{V}$, corresponding to the selection of individual tokens.
@@ 3.2 LLM-MDP
Rewards in this formulation are defined explicitly through an external verifier. Upon episode termination, a reward of 1 is given only if the terminal state $s_{H}$, representing the complete sequence generated by the LLM agent, correctly solves the given problem, otherwise, the reward is zero.
3.2.1 Structural Assumptions of the LLM-MDP
This degenerate MDP formulation adheres broadly to standard MDP properties, but introduces two distinct structural assumptions.
@@ 3.2.1 Structural Assumptions of the LLM-MDP
3.3 RL for LLM-MDP
In the context of LLM-MDP, the RL agent-environment interaction is as follows, at the beginning of each episode, the environment provides an initial state $s_{0}$, which corresponds to the question prompt or input context for a reasoning task. The LLM, acting as the RL agent, selects an action $a_{h}$ a token from the vocabulary set $\mathcal{V}$ using LLM policy $\pi_{\theta}$ based on the current state $s_{h}$, which is the sequence of previously generated tokens and the question prompt. The environment then deterministically transitions to next state $s_{h+1}$ by concatenating the selected token to the current sequence, i.e. $s_{h+1}=(s_{h},a_{h})$. This process continues until the agent generates a special end-of-sequence (EOS) token or reaches a predefined maximum sequence length H, at which point the episode terminates. Upon termination, a scalar reward is assigned by an external verifier if the final generated sequence  [^1] In practice, large language models (LLMs) often generate intermediate reasoning traces (e.g., between <Think> tags) followed by final answers (between <Sol> tags), where the <Think> section reflects chain-of-thought reasoning and the <Sol> section is evaluated by a verifier. However, as in the DeepSeek formulation, the state is defined as the sequence of tokens without distinguishing between reasoning and solution parts. We adopt the same simplification and treat the state as the just token sequence, irrespective of internal structure. constitutes a correct solution, the trajectory receives a reward of 1; otherwise, it receives a reward of 0.
In this setting, fine-tuning an LLM is equivalent to improving the agent’s policy $\pi_{\theta}$ to maximize the expected cumulative reward. RL policy gradient methods are well-suited for this task, as LLMs are transformer-based parameterized neural networks that output a probability distribution over the vocabulary given the context. In the next subsection, we provide an overview of GRPO [12], a policy gradient method used for fine-tuning LLMs.
@@ 3.4 Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) is a reinforcement learning algorithm introduced to reduce the computational cost of standard policy optimization methods, particularly in the context of fine-tuning large language models (LLMs). Traditional policy optimization methods, such as PPO, often require training a separate critic model to estimate the value function, which can be computationally expensive—especially when the critic is as large as the policy model itself. GRPO avoids this overhead by estimating the advantage function using relative rankings within a group of sampled outputs, instead of relying on a separate value network.
@@
Here, $\epsilon$ clipping factor is hyperparameter, $\hat{A}_{i,t}$ is the standardized advantage for token $o_{i,t}$, computed from the group rewards $\mathbf{r}=\{r_{1},r_{2},\dots,r_{G}\}$ as:
$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$
Additionally, a KL penalty term is included to ensure that the updated policy does not deviate too far from a reference policy, with $\beta$ controlling the strength of this regularization. The KL divergence is computed at the token level as:
@@ 4.1 Simplifying the GRPO Objective
where $\mathcal{ISR}{i,t}(\theta)$ is the importance sampling ratio:
$\mathcal{ISR}_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ (3)
Recent works ([9], [18], [19]) show that the KL penalty term has limited effect on performance, since the clipping operation already keeps the updated policy within the trust region of the old policy ([10], [11]). Therefore, we can relax the KL penalty term from the objective, simplifying it to:
@@
Now let’s assume that the importance sampling ratio $\mathcal{ISR}_{i,t}(\theta)$ are within the clipping range $(1-\varepsilon,1+\varepsilon)$ later we can put the clipping again, the objective further simplifies to:
@@
Since the relative advantage $\hat{A}{i,t}$ is computed at the output level and is constant across all tokens $o_{i,t}$ for a given response $o_{i}$, we can pull it out of the inner summation as $\hat{A}_{i,t}=\hat{A}_{i}$. This further simplifies the objective:
@@ 4.2 Decomposing by Positive and Negative Responses
Given a question $q$, we can divide its $G$ responses $\{o_{1},o_{2},...,o_{G}\}$ sampled from the old policy $\pi_{\theta_{old}}$ into the group of positive responses $\mathcal{G^{+}}$ and group of negative responses $\mathcal{G^{-}}$ based on the binary reward value assigned to each responses from the external verifier. Since every positive response of a given question $q$ the relative score will be the same we denote it as $\hat{A_{q}^{+}}$ and the relative advantage score for negative responses will also be the same we denote it as $\hat{A_{q}^{-}}$. Now we can split the objective function in positive trajectory and negative trajectory of given question $q$ and the objective equation is written as -
@@
To simplify the notation, we assume that all responses have a fixed length of $H$, with any required padding added after the EOS token. We define $A_{q}^{+}=\frac{\hat{A}_{q}^{+}}{H}$ and $A_{q}^{-}=\frac{\hat{A}_{q}^{-}}{H}$, allowing us to rewrite the objective function as:
$\mathcal{J}(\theta)$ $=\mathbb{E}\Bigg{[}\frac{1}{G}\Bigg{[}A_{q}^{+}\sum_{i=1}^{\mathcal{G^{+}}}\sum_{t=1}^{H}\mathcal{ISR}_{i,t}(\theta)+A_{q}^{-}\sum_{i=1}^{\mathcal{G^{-}}}\sum_{t=1}^{H}\mathcal{ISR}_{i,t}(\theta)\Bigg{]}\Bigg{]}$ (8)
The objective function derived above closely resembles that of filtered iterative supervised fine-tuning (F-ISFT). In standard F-ISFT, only positive responses are considered when updating the model parameters, with the goal of increasing the log-likelihood of their tokens. However, if we also incorporate negative responses into the objective—by increasing the log-likelihood for positive responses and decreasing it for negative ones—we obtain performance comparable to that of GRPO.
In this formulation, the terms $A_{q}^{+}$ and $A_{q}^{-}$ serve as adaptive weights for the positive and negative components of the objective, respectively. These relative advantage scores implicitly encode the difficulty of the question $q$. When a question is either very easy or very hard for the model, the variance among the responses tends to be low, leading GRPO to assign smaller weights to both the positive and negative terms. In contrast, for questions of moderate difficulty, where response variance is higher, GRPO assigns proportionally larger weights to both terms, thus emphasizing learning from such examples.

### v3

# LAW 2025: Bridging Language, Agent, and World Models

@@ Abstract
Reinforcement learning based post-training of large language models (LLMs) has recently gained attention, particularly following the release of DeepSeek R1 (guo2025deepseek), which applied GRPO for fine-tuning. Amid the growing hype around improved reasoning abilities attributed to RL post-training, we critically examine the formulation and assumptions underlying these methods. We start by highlighting popular structural assumptions made in modeling LLM training as an MDP, and show how they lead to a degenerate MDP that doesn’t quite need the RL/GRPO apparatus. The two critical structural assumptions include (1) making the MDP states be just a concatenation of the actions with states becoming the context window and the actions becoming the tokens in LLMs and (2) splitting the reward of a state-action trajectory uniformly across the trajectory. Our comprehensive analysis demonstrates that, due to these simplifying assumptions, the standard approach is effectively equivalent to outcome-driven supervised learning. Our experiments on benchmarks including GSM8K and Countdown using Qwen-2.5 base models and Llama-3.2 instruct models show that Filtered Iterative SFT, incorporating both positive and negative samples, achieves performance comparable to GRPO-based training. We also show that these structural assumptions indirectly incentivize RL to generate longer sequences of intermediate tokens which in turn feeds into the narrative of “RL incentivizing thinking because it generates longer thinking traces.” We continue to believe that RL writ large can be a great tool for post-training LLMs, and hope that our analysis of the limiting assumptions in the currently popular RL framework encourages work that goes beyond patching the symptoms of faulty assumptions and towards improving the underlying formulation.
@@ 1 INTRODUCTION
Recent advances in Large Language Models (LLMs) have substantially improved their ability to perform complex reasoning tasks, such as mathematical problem-solving. Typically, this is achieved through post-training techniques, including RL methods like Group Relative Policy Optimization (GRPO). These methods frame language modeling as a Markov Decision Process (MDP), where token generation is treated as a sequential decision-making task.
@@
In this work, we critically examine the foundations of applying RL to LLM post-training through the lens of the specific MDP formulation that has been popularly used to model LLMs as agents  (guo2025deepseek). Under this formulation, each state is a sequence of previously generated tokens, transitions are deterministic concatenations of tokens, rewards are assigned only at the terminal state based on external verification of correctness, and the credit assignment problem is short-circuited by effectively dividing this terminal reward across all tokens (or actions) equally. These structural assumptions sharply distinguish the setting from the general MDP framework typically considered in RL, and drive several wrong intuitions.
@@
Our contributions are twofold. Firstly, we theoretically show that under these structural assumptions, the full machinery of RL methods in LLMs post-training is somewhat redundant. By decoupling and simplifying the GRPO objective for trajectories labeled as positive (correct solutions) and negative (incorrect solutions), we demonstrate that iterative supervised fine-tuning (SFT) guided by an external verifier yields comparable performance. This equivalence implies that RL might not be imparting genuinely novel reasoning capabilities to the LLM, but rather leveraging latent capabilities inherent within the pretrained models.
@@
Secondly, we empirically demonstrate that the terminal reward distribution assumption, wherein the relative advantage score is uniformly allocated across all tokens in a response, directly contributes to the increasing response lengths observed throughout training. Contrary to the claims made in the DeepSeek-R1 (guo2025deepseek) paper, which attributes the increased response length at test-time to computational scaling, self-reflection, self-verification, and back-tracking, we show that the primary driver of response elongation is this uniform credit distribution. Since each token is incentivized equally without considering the "return-to-go" from any intermediate state, RL-based methods like GRPO inadvertently encourage verbose outputs.
This paper thus aims to clarify the misconceptions regarding the role of RL in enhancing LLM capabilities by providing a comprehensive theoretical and empirical analysis showing that idiosyncratic assumptions made in formulating the MDP influence model behavior during post training. A key implication of our analysis is that misunderstanding the consequences of these structural assumptions has led to research directions that attempt to fix superficial symptoms such as introducing length penalties (arora2025training) or selectively filtering training data (shrivastava2025sample) rather than addressing the root cause of length bias inherent in the basic formulation. We stress that our critique is directed at the specific MDP formulation currently used to model LLM post-training, rather than at RL as a general paradigm for LLMs post-training. This current, restrictive formulation does not fully realize the expressive power of RL. By highlighting these limitations, we aim to encourage the development of richer formulations that can further unlock the potential of RL in LLM post-training.
@@ 2 RELATED WORK
Supervised fine-tuning and reinforcement learning have emerged as the two main paradigms for post-training LLMs. Although early work focused heavily on SFT variants (zelikman2022star; yuan2023scaling; dong2023raft; gulcehre2023reinforced; singh2023beyond), recent efforts have shifted toward RL-inspired methods such as GRPO and its extensions (liu2025understanding; zhang2025grpo; yu2025dapo; yuan2025vapo).
@@
While the original SFT involved finetuning on the traces and solutions supplied externally, a line of work has explored improving language model reasoning via self-generated data, filtering data with verifier and iteratively fine-tuning on the filtered data, a process which we refer to as filtered iterative SFT in our work. Self-Taught Reasoner (STaR) (zelikman2022star) fine-tunes a model on its self generated data by iteratively producing rationales using greedy decoding. It retains only samples whose rationales lead to correct answers. Subsequent works have replaced greedy decoding with temperature sampling as an alternative data generation strategy. Rejection Sampling Fine-Tuning (RFT) (yuan2023scaling) collects reasoning paths by generating multiple outputs from a supervised model and filtering them using binary feedback (accepted/rejected). The model is then fine-tuned once on the accepted samples. RAFT (Reward-ranked Fine-Tuning) (dong2023raft) introduces an iterative framework consisting of three steps: (1) generate a batch of outputs from the model; (2) score the outputs using a reward function and filter high-reward samples and (3) fine-tune the model on the filtered data.
@@
Reinforced Self-Training (ReST) (gulcehre2023reinforced) similarly generates multiple output sequences per prompt, scores them with a reward function, and fine-tunes the model using a reward-weighted loss, rather than hard filtering. ReSTEM (singh2023beyond) extends this setup by fine-tuning the original base model in each iteration, instead of the latest model checkpoint. These methods collectively explore different ways to combine self-generated data, filtering, and reward signals for language model fine-tuning.
@@
The recent and growing work has focused on variations of the GRPO Framework. GRPO (shao2024deepseekmath) introduces group-relative advantages, avoiding state-based value estimation by directly assigning advantages based on relative scores within a group. Several follow-up works build on or modify GRPO in different ways. Dr-GRPO (liu2025understanding) removes standard deviation normalization and sample-length scaling from the GRPO loss, and points out that the objective may introduce a bias toward longer responses. GRPO-Lead (zhang2025grpo) extends GRPO by adding a length-dependent reward to encourage concise outputs, an explicit penalty to reduce low-precision responses, and an advantage reweighting strategy to handle limited rollout data.
@@
In parallel, DAPO (yu2025dapo) introduces a different set of techniques for long chain-of-thought settings. It uses token-level policy gradient loss and includes four key components: Clip-Higher to maintain policy diversity, Dynamic Sampling to improve training stability, Overlong Reward Shaping to reduce noise from long outputs, and the removal of flat-reward samples from training. VAPO (yuan2025vapo) builds on DAPO by introducing value model pretraining, a decoupled and length-adaptive GAE, and group sampling to better handle sparse rewards and variable-length responses.
@@
In this work, We argue that GRPO and its variants, despite their RL framing, rely on structural simplifications such as degenerate MDPs and uniform token-level reward assignment, which effectively reduce the actual learning process to iterative filtered supervised fine-tuning. A very recent work, RAFT++ (xiong2025minimalist) also extended RAFT by incorporating importance sampling and clipping and shows comparable performance as GRPO.
There have also been conflicting claims about the response length: the longer responses have both been celebrated as a sign of improved reasoning guo2025deepseek, and also viewed as a source of inefficiency, with several efforts directed at shortening them. Such efforts include sample-level loss (yu2025dapo; liu2025understanding), length-dependent penalties (yuan2025vapo; aggarwal2025l1; arora2025training), and token-based control, where Thinkless (fang2025thinkless) uses <think>/<short> tokens, while AdaptThink (zhang2025adaptthink), invokes <think> only on difficult questions. However, we show that increase in response length is more likely a side effect of the uniform advantage distribution followed with length scaling. Also, (fatemi2025concise) shows theoretical analysis that generation of longer responses stems from RL-based optimization during training.
@@ 3.1 Standard MDP Formulation
3.2 LLM-MDP and RL for LLMs
Let’s now consider the popular MDP model used for LLMs introduced in DeepSeek-R1 paper (guo2025deepseek), which extends the standard MDP Formulation by incorporating a verifier $V$ into the environment. The resulting formulation is defined as $M_{LLM}=\langle S,A,P,r,\rho,H,V\rangle$. In this scenario, the LLM acts as an agent interacting with an environment whose state space is the set of all possible token sequences up to a maximum length $H$, defined as $S=\bigcup_{k=0}^{H}\mathcal{V}^{k}.$ with $\mathcal{V}$ representing a finite vocabulary of tokens. The action space $A$ is precisely this vocabulary $\mathcal{V}$, corresponding to the selection of individual tokens.
@@ 3.2 LLM-MDP and RL for LLMs
Rewards in this formulation are defined explicitly through an external verifier. Upon episode termination, a reward of 1 is given only if the suffix of the terminal state $s_{H}$, representing the complete sequence generated by the LLM agent, contains the correct solution [^1] In practice, large language models (LLMs) often generate intermediate reasoning traces (e.g., between <Think> tags) followed by final answers (between <Sol> tags), where the <Think> section reflects chain-of-thought reasoning and the <Sol> section is evaluated by a verifier. However, as in the DeepSeek formulation, the state is defined as the sequence of tokens without distinguishing between reasoning and solution parts. We adopt the same simplification and treat the state as the just token sequence, irrespective of internal structure. to the given problem according to the external verifier; otherwise, the reward is zero.
Structural assumptions of the LLM-MDP: This LLM-MDP formulation adheres broadly to standard MDP properties, but introduces two structural assumptions.
@@
In this setting, fine-tuning an LLM is equivalent to improving the agent’s policy $\pi_{\theta}$ to maximize the expected cumulative reward. RL policy gradient methods are well-suited for this task, as LLMs are transformer-based parameterized neural networks that output a probability distribution over the vocabulary given the context. In the next subsection, we provide an overview of GRPO (shao2024deepseekmath), a policy gradient method used for fine-tuning LLMs.
@@ 3.3 Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) is a RL algorithm introduced to reduce the computational cost of standard policy optimization methods, particularly in the context of fine-tuning large language models (LLMs). Traditional policy optimization methods, such as PPO, often require training a separate critic model to estimate the value function, which can be computationally expensive especially when the critic is as large as the policy model itself. GRPO avoids this overhead by estimating the advantage function using relative rankings within a group of sampled outputs, instead of relying on a separate value network.
@@
[…]
@@ 4.1 Simplifying the GRPO Objective
where $\mathcal{ISR}_{i,t}(\theta)$ is the importance sampling ratio: $\mathcal{ISR}_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$
Recent works ((liu2025understanding), (yu2025dapo), (yuan2025vapo)) show that the KL penalty term has limited effect on performance, since the clipping operation already keeps the updated policy within the trust region of the old policy ((schulman2015trust), (schulman2017proximal)). Therefore, we can relax the KL penalty term from the objective, simplifying it to:
@@
Now let’s assume that the importance sampling ratio $\mathcal{ISR}_{i,t}(\theta)$ are within the clipping range $(1-\varepsilon,1+\varepsilon)$ (later we can put the clipping again), the objective further simplifies to:
@@
Since the relative advantage $\hat{A}_{i,t}$ is computed at the output level and is constant across all tokens $o_{i,t}$ for a given response $o_{i}$, we can pull it out of the inner summation as $\hat{A}_{i,t}=\hat{A}_{i}$. This further simplifies the objective:
@@ 4.2 Decomposing by Positive and Negative Responses
Given a question $q$, we can divide its $G$ responses $\{o_{1},o_{2},...,o_{G}\}$ sampled from the old policy $\pi_{\theta_{old}}$ into the group of positive responses $\mathcal{G^{+}}$ and group of negative responses $\mathcal{G^{-}}$ based on the binary reward value assigned to each responses from the external verifier. For every positive response of a given question $q$ the relative score will be the same and we denote it as $\hat{A_{q}^{+}}$ and the relative advantage score for negative responses will also be the same and we denote it as $\hat{A_{q}^{-}}$. Now we can split the objective function in positive trajectory and negative trajectory of given question $q$ and the objective equation is written as -
@@
We define $A_{q,i}^{+}=\frac{\hat{A}_{q}^{+}}{|o_{i}|}$ and $A_{q,i}^{-}=\frac{\hat{A}_{q}^{-}}{|o_{i}|}$, allowing us to rewrite the objective function as:
$\mathcal{J}(\theta)$ $=\mathbb{E}\Bigg[\frac{1}{G}\Bigg[\sum_{i=1}^{\mathcal{G^{+}}}A_{q,i}^{+}\sum_{t=1}^{H}\mathcal{ISR}_{i,t}(\theta)+\sum_{i=1}^{\mathcal{G^{-}}}A_{q,i}^{-}\sum_{t=1}^{H}\mathcal{ISR}_{i,t}(\theta)\Bigg]\Bigg]$ (7)
The objective function derived above closely resembles that of filtered iterative supervised fine-tuning (F-ISFT). In standard F-ISFT, only positive responses are considered when updating the model parameters, with the goal of increasing the log-likelihood of their tokens. However, if we also incorporate negative responses into the objective, by increasing the log-likelihood for positive responses and decreasing it for negative ones, we obtain performance comparable to that of GRPO.
In this formulation, the terms $A_{q,i}^{+}$ and $A_{q,i}^{-}$ serve as adaptive weights for the positive and negative components of the objective, respectively. These relative advantage scores implicitly encode the difficulty of the question $q$. When a question is either very easy or very hard for the model, the variance among the responses tends to be low, leading GRPO to assign smaller weights to both the positive and negative terms. In contrast, for questions of moderate difficulty, where response variance is higher, GRPO assigns proportionally larger weights to both terms, thus emphasizing learning from such examples.